OpenStudy (anonymous):

Solve quadratic equation. Show your work. 4x6^2= 100

OpenStudy (welshfella):

a quadratic would contain x^2 -looks like you have the question wromg

OpenStudy (anonymous):

Solve quadratic equation. Show your work. 4x^2= 100

DivineSolar (divinesolar):

List your answers if any. As i may have solved without quadratics.

OpenStudy (anonymous):

4x2 = 100 divide both sides by 4 ⦁ x2=100/4 x2=25 now you take the sqrt of both sides this cancels out the exponent of 2 on the right side. ⦁ x=sqrt 25 ⦁ 4(52)=100 ⦁ 4(25)=100 ⦁ 100=100

OpenStudy (anonymous):

I got it wrong

DivineSolar (divinesolar):

Ok let me solve it and see what i get with quadratics :)!

DivineSolar (divinesolar):

Well give me a second.

OpenStudy (anonymous):

kk

DivineSolar (divinesolar):

So your problem is 2x6^2=100? So if thats the case wouldnt a = 2, b = 6 and c = 2?

OpenStudy (anonymous):

no its 4x^2= 100

DivineSolar (divinesolar):

oh 6 was a typo?

OpenStudy (anonymous):

sorry

DivineSolar (divinesolar):

Ok so c = 100 then.

DivineSolar (divinesolar):

a = 4, b = 2.

DivineSolar (divinesolar):

Ok so here is what i get without quadratics.

OpenStudy (anonymous):

a = 4, b = 2 c = 100 is the answer

DivineSolar (divinesolar):

That is definitely not the answer, that is how you start it off.

OpenStudy (anonymous):

oh ok I got thanks

DivineSolar (divinesolar):

that is 0 quadratics though.

DivineSolar (divinesolar):

The answer i provided. IS it one of your answers? xD

OpenStudy (welshfella):

how is b=2 when theres no term in x?

DivineSolar (divinesolar):

sorry b = 0 why were u no there then.